OpenStudy (anonymous):

Explain, in complete sentences, which method you would use to solve the following system of equations and why you chose that method. Provide the solution to the system. x - 3y + 4z = -9 2x + y + 3z = 6 x - 4y + 3z = -9

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I would almost always use elimination, because it is methodical, and less confusing. write down the equations without the variables 1 -3 4 -9 2 1 3 6 1 -4 3 -9

OpenStudy (phi):

now you want to zero out the 2 just below the 1. to do that, you have to add -2 to it. to get the -2, multiply the top row by -2 and add that to the 2nd row. that becomes your new 2nd row you do not change the first row

OpenStudy (phi):

First, what is -2 times the first row ?

OpenStudy (anonymous):

-2 6 -8 18

OpenStudy (phi):

now add -2 6 -8 18 to the 2nd row -2 6 -8 18 2 1 3 6 ------------- 0 7 -5 24 that is your new 2nd row. the problem is now: 1 -3 4 -9 0 7 -5 24 1 -4 3 -9 the next step is to make the 1 in the last row zero. we must add -1. to get the -1 (using the first row), multiply the 1st row by -1 -1 * first row = -1 * 1 -3 4 -9 = -1 +3 -4 +9 add that to the 3rd row -1 +3 -4 +9 1 -4 3 -9 -------------- 0 -1 -1 0 <-- this is your new 3rd row we now have 1 -3 4 -9 0 7 -5 24 0 -1 -1 0

OpenStudy (phi):

The next step is to zero out the -1 in the bottom row 1 -3 4 -9 0 7 -5 24 0 -1 -1 0 multiply the 2nd row by 1/7 to get 0 1 -5/7 24/7 add this to the bottom row 0 1 -5/7 24/7 0 -1 -1 0 -------------- 0 0 -12/7 24/7 we now have 1 -3 4 -9 0 7 -5 24 0 0 -12/7 24/7 if you remember that there are variables x , y ,z in there, the bottom row stands for 0x + 0y -12/7 z = 24/7 or simply -12z/7 = 24/7 multiply both sides by -7/12 \[ -\frac{7}{12}\cdot -\frac{12}{7} z = \frac{24}{7} \cdot -\frac{7}{12} \] you get z= -2

OpenStudy (phi):

the middle row is short for 7y -5z= 24 but you know z= -2, so this is 7y + -5* -2 = 24 7y +10= 24 7y = 24-10 = 14 y= 14/7 =2 y=2 now use the 1st row to solve for x, using y=2 and z= -2 1 -3 4 -9 is short for x + -3*y + 4*z = -9 I assume you can find x.